# Unlock the Power of Frozen LLMs in Knowledge Graph Completion

### v2

@@ B.1 Relation Prediction
Table 8: Efficiency Comparsion between KG-LLAMA and LLAMA-MLP on the WN11 dataset. The "Procedure" column represents several key steps of these methods.
Given an incomplete triple $(h,?,t)$ as query, relation prediction aims to predict the missing relation (denoted as ????). For relation prediction, we transform the binary framework of triple classification shown in Figure 1 into a multi-class framework. The categories for classification are the relationship types within each knowledge graph. Table 3 shows our experimental results on the YAGO3-10 (Dettmers et al., 2018) dataset. We use the following prompt template, $PT_{3}$ and $PT_{4}$, to create training samples for classification, as shown in Figure 7. $PT_{4}$ introduces the $D(e)$ to generate more precise and distinguishable hidden states, linking entities in the Yago with entities in the LLM, thereby improving the overall relation prediction performance.
It is worth noting that "ALL" represents 1,079,040 training samples. LLaMA-MLP and LLAMA-DES2 (GPT) only use 0.6% of the training set to train the probing classifier. It can be seen that LLAMA-DES2 (GPT) with frozen language model parameters achieves results very close to KG-LLAMA-7B, which is fine-tuned on the full dataset.
B.2 Case Study
Table 9 shows the prediction results of different models for two samples in the WN11 test set. For the first case, it can be seen that among the methods using frozen LLMs, only LLAMA-7B-ICL, LLAMA-MLP, and LLAMA-DES-H predict correctly. However, the LLAMA-7B-ICL method predicts "True" for almost all test samples. For the second case, it can be seen that only LLAMA-MLP, LLAMA-DES-H, and KG-LLAMA-7B predict correctly. However, KG-LLAMA-7B is fine-tuned on the full training set. The correct predictions by LLAMA-MLP and LLAMA-DES-H indicate that the LLM itself has the relevant knowledge to make accurate predictions, whereas methods like LLAMA-7B-ICL and ALPACA-7B do not fully utilize the knowledge stored in the language model or generate hallucinations during the response process.
QUERY MODEL PRED LABEL
HEAD: parsnip; REL: type of; TAIL: herb; LLAMA-7B N/A TRUE
LLAMA-7B-ICL TRUE
ALPACA-7B FALSE
ALPACA-7B-ICL FALSE
LLAMA-MLP TRUE
LLAMA-DES-H TRUE
KG-LLAMA-7B TRUE
HEAD: market synset; REL: domain topic; TAIL: brain stem; LLAMA-7B N/A FALSE
LLAMA-7B-ICL TRUE
ALPACA-7B TRUE
ALPACA-7B-ICL TRUE
LLAMA-MLP FALSE
LLAMA-DES-H FALSE
KG-LLAMA-7B FALSE
Table 9: Case study (triple classification) on two representative examples selected from the WN11 dataset. "N/A" indicates that the model’s output does not contain the predicted label information (i.e., yes or no).
B.3 Efficiency Analysis
We compared the performance of the KG-LLAMA method and LLAMA-MLP on the WN11 full dataset in terms of GPU memory and time consuming metrics using an Nvidia 3090 GPU. The results are shown in Table 8. For the computation of GPU memory usage, we use "torch.cuda.memory_allocated" to obtain the peak memory usage at each step. For the calculation of GPU memory reduction factor, we focus solely on the training phase, dividing the memory usage of KG-LLAMA by that of LLAMA-MLP. For the reduction factor of time consumption, we first sum the time consumed during both the training and inference phases for each method and then divide the results. For all methods, the batch size for all procedures is set to 1. The KG-LLAMA method employs LoRA for supervised fine-tuning, and its hyperparameters are listed in Table 7, with the epoch set to 1 in the efficiency experiment. For the training stage of KG-LLAMA, we employ 8-bit quantization technique, while for the inference stage, we configure not to use beam search. For the inference stage of LLAMA-MLP, we probe both the training data and the testing data. Thus, the sum of the three times in the table corresponds to the time required for probing the training data, probing the testing data, and the MLP forward pass, respectively.
It is worth noting that in practical use, our probing method only requires a small portion of the training set to complete the training of classifiers (such as MLP, Logistic Regression, etc.). Here, we used the full dataset for fairness in testing. We can clearly observe in the Table 8 that the proposed probing method, which uses frozen language models, consumes less GPU memory at all stages compared to KG-LLAMA, even though KG-LLAMA employs parameter-efficient fine-tuning techniques like LoRA to save memory. Additionally, since our method only requires a forward pass through the LLM, it consumes significantly less time compared to KG-LLAMA.
Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
ALPACA-7B N-L 0 0.226 0.696 0.561 0.714 0.526
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
ALPACA-7B-ICL N-L 2 0.675 0.706 0.607 0.691 0.605
STRUCTURAL-AWARE-IT N-L ALL 0.907 0.959 0.752 0.927 0.881
ALPACA-DES N-L 0 0.611 - 0.590 0.548 0.525
LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
LLAMA-DES-H H-MLP ALL 0.905 - 0.742 0.937 0.866
Table 10: The experiment results (accuracy) of triple classification. ALPACA-DES combines non-generated descriptions with $PT_{2}$ for Alpaca-7B.
B.4 Additional Baselines
From Table 10, it is noteworthy that ALPACA-DES fails to outperform ALPACA-7B, suggesting that mere incorporation of entity descriptions may not yield direct performance enhancements, considering the possibility of LLMs revealing the wrong responses and being influenced by hallucinations.
Meanwhile, in our study, we did not compare our method with KOPA because it integrates KG models (such as TransE/RotatE) with language models (like LLaMA) in a hybrid approach, thereby blending multiple approaches rather than solely utilizing the capabilities of LLMs (Zhang et al., 2023). In contrast, our current research exclusively focuses on the utilization of language models (like LLaMA), rendering KOPA an unsuitable baseline for our purposes. Nonetheless, we reproduced the Structural-aware-IT model proposed by Zhang et al. (2023), which incorporates in-context learning and instruction tuning. This model samples the neighborhoods of the head entity $h$ and tail entity $t$, including the textual descriptions of neighboring triples in the demonstration prompt, followed by fine-tuning on the entire training set. Like KOPA, this approach integrates structural information from the KG with LLM capabilities. Despite this, our method, which exclusively utilizes LLMs without integrating KG structural information, achieves performance comparable to the Structural-aware-IT model, which combines both KG structure and LLM capabilities.
B.5 Analysis of the Performance in the Intermediate Layers of LLM
The selection of the optimal layer index in LLMs serves as a hyperparameter, which we may select based on empirical evidence or adjust as needed when a validation set is available. Meanwhile, our experiments found a strong consistency for the same LLM across different knowledge graphs (for example, for LLAMA, the 15th layer often yields the best KGC performance), as shown in Figure 3, which reduces the complexity of utilization.
Additionally, the superior performance of the intermediate layers of language models, as compared to those closer to the output layer, may be attributed to the potential for each layer of the model to experience hallucinations. Moreover, each layer of a language model can potentially propagate hallucinations from previous layers, manifesting as deliberate falsehoods or the expression of misleading information (Zou et al., 2023). This phenomenon can lead to scenarios where models often know the correct answer even while generating incorrect outputs (Zou et al., 2023). Concurrently, our experimental results are consistent with the conclusions of Jiang et al. (2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding. As a result, the middle layers in their experiments also tend to outperform the later layers.
B.6 Additional Analysis of the Data Efficiency
From Figure 8, we can observe the following: For the FB13 dataset, using 400 training samples, which is 0.06% of the training set, achieves 97.0% of the performance of using the entire training set; For the FB15K-237N dataset, using 0.57% of the training set achieves 99.6% of the performance of using the entire training set; For the WN18RR dataset, using 0.46% of the training set achieves 98.8% of the performance of using the entire training set; For the UMLS dataset, using 9.58% of the dataset achieves 90.6% of the performance of using the entire training set.
@@ B.6 Additional Analysis of the Data Efficiency
Figure 8: The LLAMA-DES-H prediction performance on KGC varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples, and the vertical axis represents the accuracy on the test set.
It can be observed that the LLAMA-DES-H method exhibits slightly lower data efficiency on the UMLS dataset compared to other methods. This phenomenon may be attributed to the fact that UMLS is a domain-specific knowledge graph, whereas the other datasets are encyclopedic knowledge graphs. The increased complexity of the decision boundary in the hidden states of the positive and negative example triplets generated by LLaMA necessitates a larger volume of data to adequately model the boundary.
B.7 Analysis of Methods for Entity Description Generation
From Table 2, it can be observed that the performance of LLAMA-DES2 (Tri) is significantly inferior to that of LLAMA-DES2 (GPT), and even LLAMA-MLP. This indicates that using subgraph triples as entity descriptions directly does not enhance the model’s classification performance; it may even impair its judgment. This phenomenon may be attributed to subgraph triples diverging significantly from the language model’s pre-training corpus, thus providing erroneous stimulus information. However, organizing triples into coherent semantic natural language text using methods in Appendix A.1 appears to mitigate this impact.
Figure 9: We visualize the hidden states obtained from stimulation in the FB13 test set in three-dimensional space using PCA. The high dimensional vectors are obtained using the LLAMA-DES2 (GPT) method, collecting the language model’s hidden states from the 16th layer through stimulation.
B.8 Analysis of Using Frozen LLM in KGC
It has been observed that the predictive performance achieved by the frozen language models can not achieve comparable performance with supervised fine-tuning the language models (Zhang et al., 2023; Yao et al., 2023). But the knowledge graphs used in the experiments are encyclopedic knowledge graphs, such as FB15K-237 (Toutanova and Chen, 2015), which are subsets of FB15k (Bordes et al., 2013) derived from Wikipedia. LLMs like LLaMA (Touvron et al., 2023), which are pretrained on datasets such as Wikipedia, CommonCrawl, etc., have stored this type of world knowledge (Davison et al., 2019). Liu et al. (2023) shows that prompting can effectively access the knowledge within LLMs without fine-tuning. Therefore, frozen LLMs have sufficient knowledge and capability to complete KGC tasks.

### v3

@@ 3.4. Ablation Study (For RQ2)
Table 7. Efficiency Comparsion between KG-LLAMA-7B and LLAMA-MLP on the WN11 full dataset. The ‘Procedure’ column represents several key steps of these methods.
3.5. Efficiency Study (For RQ3)
As illustrated in Figure 4, our method outperforms both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods. For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples. Additionally, combining the results from Table 1, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples. In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively. For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance. Our experimental results valid that the language model is capable of generating effective hidden states through our method, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
In addition to data efficiency, we also evaluate the computational efficiency of our method. Table 7 shows that our method achieves significantly reduced resource usage compared to parameter-efficient fine-tuning. For the computation of GPU memory usage, we use ”torch.cuda.memory_allocated” to obtain the peak memory usage at each step. The KG-LLAMA method employs LoRA and 8-bit quantization technique for supervised fine-tuning. Compared to fine-tuned 7B LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 26.11×26.$1$ overall. Notably, the training overhead of our method lies solely in the classifier, which is independent of the LLM size. Therefore, when scaling LLMs to larger sizes, such as 70B, the substantial overhead associated with backpropagation in fine-tuning becomes increasingly burdensome. In these scenarios, the advantages of our method in terms of GPU memory usage and time consumption become even more pronounced.
@@ 3.5. Efficiency Study (For RQ3)
Figure 4. The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
Figure 5. Generalization across Datasets, LLAMA-DES-H. The horizontal axis represents the training set, and the vertical axis represents the test set.
Figure 6. Performance of LLAMA-DES2 (GPT) versus the number of sampled triplets.
4. Related Work
4.1. Knowledge Graph Completion
Knowledge graphs are typically characterized by their inherent incompleteness (Hogan et al., 2021). The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural data present within knowledge graphs. For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem. KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner. And (Li et al., 2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning. More recently, KG-LLAMA (Yao et al., 2024) is the first to adopt a decoder-only LLM architecture for KGC by formulating the task as question answering through an instruction-tuning-based prompt, followed by parameter-efficient fine-tuning on LLaMA. Notably, KG-LLAMA leverages only LLaMA’s own capabilities without incorporating additional hybrid approaches, which aligns with the core principle of our proposed framework.
4.2. Large Language Models
Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content. As highlighted by (Saier et al., 2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters. This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures. (Ramrakhiyani et al., 2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques. Similarly, (Jin et al., 2024) used probing to measure the abilities of different LLM layers to learn various levels of concepts. In addition, (Liu et al., 2024) investigated whether a universal truthfulness hyperplane exists within LLMs by employing probing methods.
5. Conclusion
This paper proposes a novel framework that leverages the latent representation space of LLMs to improve KGC. By integrating subgraph-based entity descriptions, our approach bridges the semantic gap between LLMs and KGs, enabling LLMs to achieve performance on par with their fine-tuned counterparts. Simultaneously, our method enhances data efficiency and saves a significant amount of GPU memory and time compared to parameter-efficient fine-tuning. This work explores the potential of LLMs in KGC tasks with both efficacy and efficiency. Compared to fine-tuning, we hit the core of KGC problem in a better way.